# How do Large Language Models Handle Multilingualism?

### v1

@@ Abstract
Large language models (LLMs) demonstrate remarkable performance across a spectrum of languages. In this work, we delve into the question: How do LLMs handle multilingualism? We introduce a framework that depicts LLMs’ processing of multilingual inputs: In the first several layers, LLMs understand the question, converting multilingual inputs into English to facilitate the task-solving phase. In the intermediate layers, LLMs engage in problem-solving by thinking in English and incorporating multilingual knowledge to obtain factual content, leveraging the self-attention and feed-forward structures, respectively. In the last several layers, LLMs generate responses that align with the original language of the query. In addition, we investigate the existence of language-specific neurons when processing a certain language. To detect neurons activated by the input language, even without labels, we innovatively design a Parallel Language specific Neuron Detection (PLND) method that effectively measures the significance of neurons when handling multilingual inputs. By comprehensive ablation analysis through deactivating neurons of different layers and structures, we verify the framework that we propose. Additionally, we demonstrate that we can utilize such a framework to effectively enhance the multilingual ability with much less training effort. [^1] Our code implementation will be publicly available at https://github.com/DAMO-NLP-SG/multilingual_analysis
@@ 1 Introduction
Recent advancements in large language models (LLMs) such as PaLM (Chowdhery et al., 2022), GPT-4 (OpenAI, 2023), LLaMA (Touvron et al., 2023), and Mistral (Jiang et al., 2023) have dramatically transformed traditional natural language processing (NLP) tasks, integrating seamlessly into daily and professional uses. Thanks to their extensive pretraining on massive corpora mixed with different languages, these models demonstrate remarkable capabilities in understanding and generating text across multiple languages (Huang et al., 2023; Zhu et al., 2023; Zhang et al., 2023; Zhao et al., 2024). Despite their proven effectiveness, the intricate workings of their multilingual processing mechanisms remain largely unclear, which leads to an important research question: How do large language models handle multilingualism?
Some studies explore the multilingual capabilities of language models, focusing on their cross-lingual performance or structural commonalities between languages K et al. (2020); Doddapaneni et al. (2021); Chai et al. (2022). More recently, to understand the relation of certain abilities with specific model architecture especially the dominant Transformer (Vaswani et al., 2017) architecture, studies such as Hou et al. (2023); Stolfo et al. (2023); Friedman et al. (2023) explore the reasoning abilities of LLMs with self-attention layers. While others like Geva et al. (2021); Dai et al. (2022a); Meng et al. (2022) focus on the feed-forward layers, regarding them as key-value memories for storing factual knowledge. However, these works often examine components in isolation and lack a comprehensive framework that encompasses the entire end-to-end process. More importantly, the working mechanism of multilingualism is largely unaddressed in these interpretations.
@@
Figure 1: Distribution of English and non-English tokens among layers given non-English instructions.
In this study, to gain an initial understanding of the multilingual mechanisms within LLMs, we analyze the decoded embeddings after each layer when processing inputs in various languages other than English. We then classify these embeddings as corresponding to either English or non-English tokens (see details in Appendix A). As illustrated in Figure 1, the models initially represent non-English user instructions in non-English form. However, as the instructions are processed through the layers of the model, the representation surprisingly becomes English-centric. In the final layers, we observe a reversion to a predominance of non-English embeddings, as the initial instructions are posted in non-English languages. This pattern suggests a complex interplay between languages within LLMs, as they negotiate between instruction fidelity and a possible English-oriented processing bias.
Figure 2: Our proposed multilingual workflow of LLMs.
Motivated by these observations, we propose a new framework shown in Figure 2 that conceptualizes the operational stages of LLMs when processing multilingual inputs: In the first several layers, LLMs understand the user input and convert the diverse linguistic features into a unified representation. Transitioning into the task-solving phase, LLMs engage in solving the tasks by thinking in English and incorporating multilingual knowledge to obtain factual contents, leveraging the self-attention and feed-forward structures, respectively. Finally, models generate responses that align with the original language of the query. Such a framework echoes the observations we have above and also some findings in previous studies, and thus reveals the LLMs’ ability to shift among languages, allowing them to comprehend, reason, and respond effectively to multilingual tasks.
The subsequent question that arises is whether LLMs engage all available neurons of certain components or only a particular subset when processing a single language, which essentially investigates the presence of “language-specific neurons”. To tackle this question, it is crucial to pinpoint which neurons are activated by the input, even without explicit labels of certain tasks. Therefore, we further develop a novel approach called Parallel Language-specific Neuron Detection (PLND), which effectively measures the significance of individual neurons in relation to the input provided. With PLND, we can identify language-specific neurons by feeding free text corpus of that language into the model and isolating neurons that consistently activate in responses. Experimental results show that by deactivating language-specific neurons which account for only $0.13\%$ of all neurons, LLMs’ performance on a summarization task could drop by $99\%$.
With the proposed PLND method, we then conduct extensive experiments to validate our hypothesis illustrated in Figure 2. We selectively disable neuron groups to study their effect on LLM multilingualism. Notably, deactivating the language-specific neurons in the understanding layer leaves English XQuAD performance stable but decreases non-English performance by $14\%$. Other tasks including reasoning, knowledge question answering and NLG also exhibit the same characteristics. Moreover, enhancing the multilingual capabilities of LLMs can be achieved by fine-tuning language-specific neurons with merely $200$ contextual examples. This method results in a substantial increase in model performance, with a $7.4\%$ relative improvement on the XQuAD and an $8.9\%$ enhancement on XLSum, reducing the size of the required training dataset significantly.
@@ 2 Parallel Language-specific Neuron Detection (PLND)
In this section, we introduce a neuron detection method called PLND for investigating the presence of language-specific neurons.
@@ 2.1 Sequential Neuron Detection
To identify neurons responsible for a specific language, it is crucial to discern the significance of a neuron in relation to the inference of a given input. We denote the input of $i$-th layer in Transformer (Vaswani et al., 2017) as $h_{i}$, with the corresponding output represented as $h_{i+1}=T_{i}(h_{i})$, where $T_{i}$ represents the parameters of the $i$-th layer. For a specific neuron, denoted as $N^{(i)}_{k}$, within the $i$-th layer—whether located in the attention or feed-forward layer—the importance is quantified as the difference between output when $N^{(i)}_{k}$ is either activated or deactivated. Formally, it is defined as
$\text{Imp}(N^{(i)}_{k}|h_{i})=\|T_{i}\backslash N^{(i)}_{k}(h_{i})-T_{i}(h_{i})\|_{2},$ (1)
where $T_{i}\backslash N_{k}(\cdot)$ denotes deactivating $N^{(i)}_{k}$ in $T_{i}$. Then, with a set of the corpus in the specific language, denoted as $\mathcal{C}=\{c_{1},\cdots,c_{l},\cdots,c_{n}\}$, we can calculate the importance of each neuron in each layer to each corpus. Furthermore, we can select neurons that are important to all corpus in $\mathcal{C}$, i.e.,
$\text{Imp}(N^{(i)}_{k}|c_{l})\geq\epsilon,\;\forall c_{l}\in\mathcal{C},$ (2)
where $\epsilon$ is the pre-defined threshold. However, it is super time-consuming to traverse all neurons and all inputs sequentially. Therefore, we need to design a parallel algorithm for acceleration.
@@ 2.2 Parallel Neuron Detection
Feed-Forward Layer
In Llama2 (Touvron et al., 2023), the $\text{FFN}(x)$ is defined as
$\Big{(}\text{SiLU}\big{(}W_{gate}(x)\big{)}\cdot W_{up}(x)\Big{)}W_{down},$ (3)
where ${x\in\mathbb{R}^{l\times d_{model}}}$, ${W_{gate}\in\mathbb{R}^{d_{model}\times d_{inter}}}$, ${W_{down}\in\mathbb{R}^{d_{inter}\times d_{model}}}$. We denote hidden embedding before $W_{down}$ as $h_{\text{ffn}}$. When deactivating the $k$-th neuron of $W_{up}$,
$\text{Imp}(W_{up}[:,k]|x)=\|\hat{\text{FFN}}(x)-\text{FFN}(x)\|_{2}$ (4)
$=\Big{\|}\big{(}h_{\text{ffn}}\cdot\texttt{Mask}[k]\big{)}W_{down}(x)\Big{\|}_{2},$
@@
where $\texttt{Mask}[k]$ is a vector of length $d_{inter}$ with the $k$-th element as $1$ and others as 00. For calculating $\text{Imp}(W_{up}[:,k]|x)$ for all neurons in $W_{up}$ parallelly, we introduce a diagonal mask matrix of size $(d_{inter},d_{inter})$, denoted as Mask. Therefore,
$\text{Imp}(W_{up}|x)=\|(h_{\text{ffn}}\cdot\texttt{Mask})W_{down}(x)\|_{2}.$ (5)
@@
Furthermore, we find that deactivating the $k$-th neuron of $W_{down}$ is equivalent to deactivating the $k$-th neuron in $W_{up}$ as they all set $h_{\text{ffn}}[k]=0$. Therefore $\text{Imp}(W_{down}|x)$ can be obtain by Equation (5).
Self-Attention Layer
@@
For the input $x$ of length $l$, the self-attention layer is defined as
$\text{Softmax}\big{(}\frac{W_{Q}(x)W_{K}^{T}(x)}{\sqrt{d}}\big{)}W_{V}(x),$ (6)
@@
where $W_{Q}\in\mathbb{R}^{d_{model}\times d_{mid}}$, $W_{K}\in\mathbb{R}^{d_{model}\times d_{mid}}$, $W_{V}\in\mathbb{R}^{d_{model}\times d_{mid}}$. [^2] In Vicuna and Mistral, $d_{model}=d_{mid}$, but we use different notations to avoid ambiguity. As $W_{V}(x)$ is a linear layer, $\text{Imp}(W_{V}|x)$ can be obtained following Equation (5). In the case of $W_{Q}$, when deactivating the $k$-th neuron, $\hat{W}_{Q}\leftarrow W_{Q}[:,k]=0$, we aim to obtain $\text{Imp}(W_{Q}[:,k]|x)$. Firstly, we calculate the difference in attention weight, i.e., $W_{Q}(x)W_{K}^{T}(x)$.
$\Delta_{k}$ $=\hat{W}_{Q}(x)W_{K}^{T}(x)-W_{Q}(x)W_{K}^{T}(x)$ (7)
$=W_{Q}(x)[:,k]W_{K}(x)[k,:]\in\mathbb{R}^{l\times l}$
Then, the importance of $W_{Q}[:,k]$ can be defined as
$\text{Imp}(W_{Q}[k,:]|x)$ (8)
$\approx\|\hat{\text{attention}}(x)-\text{attention}(x)\|_{2}$
$\approx\Big{\|}\text{softmax}\big{(}\frac{W_{Q}(x)W_{K}^{T}(x)-\Delta_{k}}{\sqrt{d}}\big{)}-$
$\quad\;\;\text{softmax}\big{(}\frac{W_{Q}(x)W_{K}^{T}(x)}{\sqrt{d}}\big{)}\Big{\|}_{2}$
This process can also be calculated parallelly, i.e.,
$\Delta=$ $\;\hat{W}_{Q}(x)W_{K}^{T}(x)-W_{Q}(x)W_{K}^{T}(x)$ (9)
$=$ $\;W_{Q}(x).resize(l,1,d_{mid})\times$
$\;W_{K}(x).resize(1,l,d_{mid})\in\mathbb{R}^{l\times l\times d_{mid}}$
Then, the importance of $W_{Q}$ can be defined as
$\text{Imp}(W_{Q}|x)$ $\approx\Big{\|}\text{softmax}\big{(}\frac{W_{Q}(x)W_{K}^{T}(x)-\Delta}{\sqrt{d}}\big{)}-$
$\quad\;\;\text{softmax}\big{(}\frac{W_{Q}(x)W_{K}^{T}(x)}{\sqrt{d}}\big{)}\Big{\|}_{2}.$
$\text{Imp}(W_{K}|x)$ can be calculated the same way.
3 Investigate Language-Specific Neurons
@@
In this section, we apply the PLND method to selected languages and models in order to confirm the existence of language-specific neurons and investigate the relationships between languages.
3.1 Experimental Setup
Models
We test two open-source models that perform well on multilingual tasks, including Vicuna-7b-v1.5 (Chiang et al., 2023) and Mistral-7b-v1.0 (Jiang et al., 2023). Vicuna-7b-v1.5 is an open-source chat model fine-tuned on top of LlaMa 2 (Touvron et al., 2023) via instruction fine-tuning and has shown good multilingual instruction following ability. [^3] We do not directly utilize Llama2-chat as it does not follow multilingual instructions, consistently responding in English regardless of the language of the query. Mistral-7b-v1.0 is fine-tuned on Mistral-7b. For simplicity, we abbreviate them as Vicuna and Mistral hereafter to represent the two models respectively.
Languages
We adopt $6$ languages including English (En), German (De), French (Fr), Chinese (Zh), Spanish (Es), and Russian (Ru). These languages were selected because their training corpora represent over $0.1\%$ of the total pre-training data for Llama2. Furthermore, both Vicuna and Mistral are designed to tackle problems across these languages effectively, and their performance in these languages is much higher than in other low-resource languages such as Thai, Swahili, and Greek.
Corpus
To compile a language-specific corpus without task-specific considerations, we utilize OSCAR (Caswell et al., 2020) corpus which contains web crawling texts for each language. Our selection criterion for the number of contexts is based on achieving substantial coverage of each language’s vocabulary, ensuring that the selected contexts provide a representative sample of the language. Details are shown in Table 1, where “corpus size” indicates the number of contexts selected, “corpus vocab” represents the vocabulary coverage within the selected contexts, “vocab size” refers to the number of vocabularies of that language.
Language En De Fr Zh Es Ru
Corpus Size $180$k $30$k $50$k $20$k $20$k $20$k
Corpus Vocab $249$k $154$k $134$k $198$k $90$k $144$k
Vocab Size $273$k $148$k $135$k $329$k $93$k $150$k
Table 1: Corpus details across languages are tailored to encompass the majority of each language’s vocabulary.
3.2 Existence of Language-Specific Neurons
@@
Using PLND, we feed a corpus in a specific language to LLMs and identify the neurons that are consistently activated. These language-specific neurons are responsible for processing queries in that language. To ascertain whether these neurons are genuinely language-specific, we assess the performance of LLMs in corresponding languages when these neurons are deactivated versus when the same number of randomly sampled neurons is deactivated. We utilize the summarization task with XLSum dataset (Hasan et al., 2021) to showcase the capabilities of LLMs, aiming to demonstrate that the resulting output can be devoid of meaning.
Method Fr Zh Es Ru Avg.
@@
Deact-Rand. $14.1$ $61.6$ $10.4$ $20.8$ $26.7$
Deact-Lang. $\mathbf{0.83}$ $\mathbf{0.00}$ $\mathbf{0.24}$ $\mathbf{0.42}$ $\mathbf{0.37}$
@@
Deact-Rand. $15.4$ $55.9$ $10.2$ $21.2$ $25.7$
Deact-Lang. $\mathbf{0.21}$ $\mathbf{0.39}$ $\mathbf{0.15}$ $\mathbf{0.07}$ $\mathbf{0.21}$
Table 2: Multilingual performance of LLMs upon deactivating neurons, where “Lang.” denotes language-specific neurons and “Rand.” indicates an equivalent number of neurons selected at random.
Table 2 demonstrates the decline of multilingual capabilities when deactivating language-specific neurons. Although just deactivating around $0.13\%$ neurons, LLMs lose multilingual capabilities including understanding multilingual inputs, handling multilingual questions, and generating multilingual outputs. In contrast, deactivating the same number of randomly selected neurons does not yield any difference. Therefore, the detected neurons are language-specific and related to handling corresponding multilingual inputs.
3.3 Interrelation of Language-Specific Neurons Across Languages
Using neurons identified by PLND, we investigate the relationships of two languages via the degree of overlap between their language-specific neurons:
We can observe that in both Mistral and Vicuna, the intersection with English from other languages is relatively limited, suggesting that English possesses a predominant number of language-specific neurons. Additionally, there is a pronounced tendency for languages belonging to the same family to demonstrate a higher degree of overlap with each other. Moreover, the feed-forward structure typically exhibits a higher degree of consistency in overlap across various languages, due to the shared world knowledge embedded within the neurons that is accessible to multiple languages.
4 How do LLMs Handle Multilingualism?
In this section, we validate the workflow presented in Figure 2 by systematically deactivating language-specific neurons across different layers and components to examine their impact on a range of tasks.
4.1 Experimental Setup

### v2

@@ Abstract
Large language models (LLMs) have demonstrated impressive capabilities across diverse languages. This study explores how LLMs handle multilingualism. Based on observed language ratio shifts among layers and the relationships between network structures and certain capabilities, we hypothesize the LLM’s multilingual workflow (MWork): LLMs initially understand the query, converting multilingual inputs into English for task-solving. In the intermediate layers, they employ English for thinking and incorporate multilingual knowledge with self-attention and feed-forward structures, respectively. In the final layers, LLMs generate responses aligned with the original language of the query. To verify MWork, we introduce Parallel Language-specific Neuron Detection (PLND) to identify activated neurons for inputs in different languages without any labeled data. Using PLND, we validate MWork through extensive experiments involving the deactivation of language-specific neurons across various layers and structures. Moreover, MWork allows fine-tuning of language-specific neurons with a small dataset, enhancing multilingual abilities in a specific language without compromising others. This approach results in an average improvement of $3.6\%$ for high-resource languages and $2.3\%$ for low-resource languages across all tasks with just $400$ documents.
@@ 1 Introduction
Recent advancements in large language models (LLMs)  (OpenAI, 2023; Touvron et al., 2023; Team et al., 2023) have dramatically transformed the field of natural language processing (NLP). Thanks to the extensive pretraining on massive corpora mixed with different languages, these models demonstrate remarkable capabilities in understanding and generating text across multiple languages (Huang et al., 2023; Zhu et al., 2023; Zhang et al., 2023a; Zhao et al., 2024a). Despite these advancements, the intricate mechanism of their multilingual processing behavior remains largely unclear, which leads to an important research question: How do large language models handle multilingualism?
To understand the working mechanism of LLMs, existing studies mainly focus on the relationship between model architectures and certain capabilities, with some investigating reasoning abilities with self-attention layers (Hou et al., 2023; Stolfo et al., 2023; Friedman et al., 2023), and others interpreting feed-forward layers as key-value memories for storing factual knowledge (Geva et al., 2021; Dai et al., 2022a; Meng et al., 2022). However, these works solely center on English and neglect the multilingual features of LLMs in their interpretations.
@@
Figure 1: Ratio of English and non-English tokens among layers given non-English queries.
To gain an initial understanding of the multilingual mechanism of LLMs, we test LLMs with various non-English queries and decode the hidden embeddings of each layer to tokens within the LLM’s vocabulary. Subsequently, we classify these decoded tokens into either English or non-English, and analyze the ratio. Figure 1 illustrates the ratio of English and non-English tokens for each layer of two LLMs. We observe that non-English queries initially generate non-English embeddings as expected. However, as queries progress through the middle layers, the representations become English-centric. In the final layers, there is a reversion to predominantly non-English embeddings, matching the non-English queries.
Figure 2: Our hypothesized multilingual workflow MWork.
Motivated by the observed transformation above, we hypothesize a three-stage multilingual workflow: understanding, task-solving, and generating. This involves understanding the original non-English queries and interpreting them in English, solving tasks in English, and reverting outputs back to the original language. Furthermore, building upon previous studies that link self-attention structures to reasoning and feed-forward structures to factual knowledge storage (Hou et al., 2023; Geva et al., 2021), we further decouple the task-solving stage into reasoning with self-attention structures and extracting multilingual knowledge with feed-forward structures. Therefore, our hypothesized Multilingual Workflow (MWork) illustrated in Figure 2 outlines the three operational stages of LLMs in processing multilingual queries: Initially, LLMs understand queries by converting diverse linguistic features into a unified representation. In the task-solving phase, LLMs think in English and incorporate multilingual knowledge to obtain factual content, using self-attention and feed-forward structures, respectively. Finally, models generate responses in the original language as the original query.
To verify the proposed MWork workflow, we extract language-specific parameters and selectively deactivate them within different structures, thereby assessing the functionality of corresponding structures and validating our hypothesis. To identify the parameters to be activated, we develop a novel approach called Parallel Language-specific Neuron Detection (PLND). Unlike existing methods that rely on fine-tuning(Frankle and Carbin, 2018; Aghajanyan et al., 2021; Zhang et al., 2023b), labeled data (Tang et al., 2024; Liu et al., 2024), or parallel corpora (Libovickỳ et al., 2020; Tanti et al., 2021; Zhang et al., 2024) to detect activated parameters, PLND measures the significance of individual neurons with respect to the input in both attention and feed-forward structures without any labeled data or parameter adjustments. Using PLND, we identify language-specific neurons with free text corpus of that language and isolate consistently activated neurons. We find that by deactivating language-specific neurons which account for only $0.13\%$ of all neurons, LLMs’ performance on a summarization task could drop by $99\%$.
We then extensively verify the hypothesized MWork framework using the proposed PLND method. Employing various benchmark tasks, including XQuAD (Artetxe et al., 2020) for understanding, MGSM (Shi et al., 2022) for reasoning, X-CSQA (Lin et al., 2021) for knowledge extraction, and XLSum for generation (Hasan et al., 2021), we selectively deactivate language-specific neurons in each component and verify the functionality of the component by observing a significant decline in performance on the corresponding task. For example, when deactivating the language-specific neurons in the understanding layer, the performance on the multilingual understanding task XQuAD remains stable in English, while experiencing a decrease of $14\%$ in non-English languages. Other tasks exhibit the same characteristics when deactivating corresponding neurons. More importantly, with the verified MWork framework, enhancing the multilingual capabilities of LLMs can thus be achieved through the fine-tuning of language-specific neurons for certain capabilities. With a remarkable reduction in the training corpus size to a mere few hundred documents, this fine-tuning procedure enhances the multilingual capabilities of LLMs for both high-resource and low-resource languages by an average of $3.6\%$ and $2.3\%$ across all tasks, respectively. Notably, even without an English training corpus, there is a noticeable improvement in English performance, as the enhancement of language-specific neurons yields greater accuracy in enhancing specific languages, while simultaneously ensuring a clear division of parameters among different languages. In summary, the verified MWork reveals how LLMs handle multilingual tasks and offers an effective approach for conducting language-specific enhancements without compromising performance in other languages.
@@ 2 Parallel Language-specific Neuron Detection (PLND)
To verify the hypothesized workflow, we propose PLND that effectively detects language-specific neurons without relying on any labeled data.
@@ 2.1 Sequential Neuron Detection
We define a neuron as a single row or column of a parameter matrix of a language model. To identify neurons responsible for a specific language, it is crucial to discern the significance of a neuron with respect to the inference of a given input. Specifically, when processing the input $c$ in the model, we denote the hidden embedding before the $i$-th layer in Transformer (Vaswani et al., 2017) as $h_{i}$, and the hidden embedding after the $i$-th layer as $h_{i+1}=T_{i}(h_{i})$, where $T_{i}$ represents the parameters of the $i$-th layer. For a specific neuron within the $i$-th layer, denoted as $N^{(i)}$, either located in the attention or feed-forward network, we quantify its importance in processing the input $c$ by measuring the difference in the hidden embedding after the $i$-th layer, i.e., $h_{i+1}$, when $N^{(i)}$ is activated or deactivated. Formally, the impact of neuron $N^{(i)}$ for input $c$ is defined as
$\text{Imp}(N^{(i)}|c)=\|T_{i}\backslash N^{(i)}(h_{i})-T_{i}(h_{i})\|_{2},$ (1)
where $T_{i}\backslash N^{(i)}(\cdot)$ denotes deactivating $N^{(i)}$ in $T_{i}$, i.e., setting all parameters of the neuron $N^{(i)}$ to zero. With a set of $n$ corpus in a specific language, denoted as $\mathcal{C}=\{c_{1},\cdots,c_{l},\cdots,c_{n}\}$, we calculate the importance of each neuron in each layer to each corpus. Furthermore, we can obtain language-specific neurons that are important to all corpus in that language, i.e.,
$\{N^{(i)}\mid\text{Imp}(N^{(i)}|c_{l})\geq\epsilon,\forall c_{l}\in\mathcal{C}\},$ (2)
where $\epsilon$ is the pre-defined threshold.
@@ 2.2 Parallel Neuron Detection
The sequential neuron detection is time-consuming, requiring traversal of all neurons and inputs sequentially. To address this, we further propose a parallel algorithm for accelerating the process.
Feed-Forward Network (FFN)
In the latest open-source models, when processing input $c$, the feed-forward network in a certain layer is defined as
$\text{FFN}(x)=\Big{(}\text{SiLU}\big{(}W_{gate}(x)\big{)}\cdot W_{up}(x)\Big{)}W_{down},$ (3)
[…]
$\text{Imp}(W_{up}[:,k]|c)=\|\hat{\text{FFN}}(x)-\text{FFN}(x)\|_{2}=\Big{\|}\big{(}h_{\text{ffn}}\cdot\texttt{Mask}[k]\big{)}W_{down}(x)\Big{\|}_{2},$ (4)
@@
where $h_{\text{ffn}}\in d_{inter}$ represents the embedding before $W_{down}$, and $\texttt{Mask}[k]\in d_{inter}$ is a vector with the $k$-th element equal to $1$ and the rest equal to 00. To calculate $\text{Imp}(W_{up}[:,k]|c)$ for $k\in d_{inter}$ parallelly, we introduce a diagonal mask matrix of size $(d_{inter},d_{inter})$, denoted as Mask. Therefore,
$\text{Imp}(W_{up}|c)=\|(h_{\text{ffn}}\cdot\texttt{Mask})W_{down}(x)\|_{2}.$ (5)
@@
Furthermore, we observe that deactivating the $k$-th neuron of $W_{down}$ is equivalent to deactivating the $k$-th neuron in $W_{up}$, as they both result in $h_{\text{ffn}}[k]=0$. Hence, we can also derive $\text{Imp}(W_{down}|c)$ by employing Equation (5).
Self-Attention Network
@@
When processing input $c$, the self-attention network in a certain layer is
$\text{Attention}(x)=\text{Softmax}\big{(}\frac{W_{Q}(x)W_{K}^{T}(x)}{\sqrt{d}}\big{)}W_{V}(x),$ (6)
@@
where $W_{Q},W_{K},W_{V}\in\mathbb{R}^{d_{model}\times d_{mid}}$.  [^2] In some models like Vicuna and Mistral, $d_{model}=d_{mid}$, but we use different notations to avoid ambiguity. Since $W_{V}(x)$ is not in the non-linear softmax calculation, we can calculate $\text{Imp}(W_{V}|c)$ by applying Equation (5). For $W_{Q}$, we obtain $\text{Imp}(W_{Q}[:,k]|c)$ by deactivating its $k$-th neuron, specifically, $\hat{W}_{Q}\leftarrow W_{Q}[:,k]=0$. Firstly, we calculate the difference in attention weight before and after deactivation, prior to scaling and softmax,
$\Delta_{k}(x)$ $=\hat{W}_{Q}(x)W_{K}^{T}(x)-W_{Q}(x)W_{K}^{T}(x)=W_{Q}(x)[:,k]W_{K}(x)[k,:]\in\mathbb{R}^{l\times l}.$ (7)
Next, as the changes in attention exhibit a positive correlation with the changes in the output of this layer, the importance of $W_{Q}[:,k]$ in processing $c$, as defined in Equation 1, can be approximated as
$\text{Imp}(W_{Q}[:,k]|c)$ $\approx\|\hat{\text{attention}}(x)-\text{attention}(x)\|_{2}$ (8)
$\approx\Big{\|}\text{softmax}\big{(}\frac{W_{Q}(x)W_{K}^{T}(x)-\Delta_{k}(x)}{\sqrt{d}}\big{)}-\text{softmax}\big{(}\frac{W_{Q}(x)W_{K}^{T}(x)}{\sqrt{d}}\big{)}\Big{\|}_{2}.$
This process can also be calculated in parallel, specifically,
$\Delta(x)=$ $\;\hat{W}_{Q}(x)W_{K}^{T}(x)-W_{Q}(x)W_{K}^{T}(x)$ (9)
$=$ $\;W_{Q}(x).resize(l,1,d_{mid})\times W_{K}(x).resize(1,l,d_{mid})\in\mathbb{R}^{l\times l\times d_{mid}}.$
Therefore, the importance of $W_{Q}$ in processing input $c$ is calculated by
$\text{Imp}(W_{Q}|c)$ $\approx\Big{\|}\text{softmax}\big{(}\frac{W_{Q}(x)W_{K}^{T}(x)-\Delta(x)}{\sqrt{d}}\big{)}-\text{softmax}\big{(}\frac{W_{Q}(x)W_{K}^{T}(x)}{\sqrt{d}}\big{)}\Big{\|}_{2}.$ (10)
Similarly, since $W_{K}$ is symmetrical to $W_{Q}$, $\text{Imp}(W_{K}|c)$ can be calculated in the same way.
2.3 Detection of Language-Specific Neurons
@@
We then apply PLND to selected languages and models to validate its effectiveness in detecting language-specific neurons and to further investigate the relationships between languages.
Experimental Setup.
We test two open-source models that perform well on multilingual tasks, including Vicuna-7b-v1.5333We do not directly utilize Llama2-chat as it does not follow multilingual instructions, consistently responding in English regardless of the language of the query. (Chiang et al., 2023) and Mistral-7b-Instruct-v0.2 (Jiang et al., 2023). For simplicity, we abbreviate them as Vicuna and Mistral hereafter to represent the two models respectively. We select the text summarization task with the XLSum (Hasan et al., 2021) dataset as the reference task to evaluate multilingual performance as it requires the model to comprehend the input text and generate a coherent fragment. We adopt $4$ high-resource languages including French (Fr), Chinese (Zh), Spanish (Es), and Russian (Ru), as their initial performance on those languages is already quite reasonable for observing the multilingual processing mechanism. Furthermore, we utilize OSCAR (Caswell et al., 2020) corpus which contains web crawling texts for each language to compile a language-specific corpus without task-specific considerations, and details are presented in Appendix B.
Existence of Language-Specific Neurons
@@
Using PLND, we feed a corpus in a specific language to LLMs and identify neurons that are consistently activated, which are responsible for processing queries in that language. To ascertain whether these neurons are genuinely language-specific, we assess the performance of LLMs in corresponding languages when these neurons are deactivated versus when the same number of randomly sampled neurons are deactivated.
Table 1: Multilingual performance on XLSum when deactivating language-specific neurons (“Lang-Spec”) and an equivalent number of randomly selected neurons (“Random”).
Model Method Fr Zh Es Ru Avg.
@@
Deactivate Random $14.1$ $61.6$ $10.4$ $20.8$ $26.7$
Deactivate Lang-Spec $\mathbf{0.83}$ $\mathbf{0.00}$ $\mathbf{0.24}$ $\mathbf{0.42}$ $\mathbf{0.37}$
@@
Deactivate Random $15.4$ $55.9$ $10.2$ $21.2$ $25.7$
Deactivate Lang-Spec $\mathbf{0.21}$ $\mathbf{0.39}$ $\mathbf{0.15}$ $\mathbf{0.07}$ $\mathbf{0.21}$
Table 1 demonstrates the decline of multilingual capabilities when deactivating language-specific neurons. Although just deactivating around $0.13\%$ neurons, LLMs lose their multilingual capabilities and fail to generate meaningful content. In contrast, deactivating the same number of randomly selected neurons does not yield any difference. Therefore, the detected neurons are language-specific and related to handling corresponding multilingual inputs. Furthermore, we investigate the degree of overlap among their language-specific neurons in Appendix C.
3 Multilingual Workflow (MWork) of LLMs
3.1 MWork
Figure 3: Number of language-specific neurons when processing multilingual queries.
By classifying the hidden representations of each layer in LLMs into English or non-English (as shown in Figure 1), we can observe the shift from non-English to English-centric, and back to non-English with the progression through the layers. This motivates us to hypothesize a three-stage multilingual workflow: understanding the original non-English queries and interpreting them in English, task-solving in English, and generating back to the original language. Nevertheless, the presence of certain non-English tokens during the English-centric task-solving stage inspires us to further investigate this stage.
With the proposed PLND method, we extract language-specific neurons from attention and feed-forward structures when processing various multilingual queries. We plot the average number of activated language-specific neurons when processing each query in Figure 3. Notably, the number of language-specific neurons decreases within the self-attention structure in the task-solving layer but remains consistent across the layers of the feed-forward structure. This decline implies a reliance on the English language for thinking while extracting multilingual knowledge to support query processing, which is also consistent with (Geva et al., 2021)’s interpretation of the feed-forward structure as key-value memories for knowledge extraction. Therefore, we further decompose the task-solving layer into two parts: thinking in English and extracting knowledge in a multilingual context.
Considering the above insights, we propose the MWork hypothesis for explaining LLM’s multilingual workflow: LLMs first understand user input by unifying diverse linguistic features. They then engage in the task-solving phase, employing English for thinking and leveraging multilingual knowledge through self-attention and feed-forward structures, respectively. Finally, the models generate responses aligned with the query’s original language.
3.2 Verification Experiment Setup
To verify MWork, we selectively deactivate language-specific neurons from each component. Then its functionality can be verified if this deactivation results in minimal impact on English performance while exhibiting a notable decline in multilingual performance for the corresponding task.